OpenStudy (castiel):

If sin(x-(3pi/2))=-4/5, pi/2

OpenStudy (cwrw238):

as x is in the second quadrant its sine is 4/5 sin x = 2 sin(x/2)cos(x/2) so sin(x/2)*cos(x/2) = sin x / 2 = 2/5

OpenStudy (castiel):

Oh man thank you, here I was making up formulas, I completely forgot sinx=2sin(x/2)cos(x/2). Just a small correction that sinx=3/5 so the final result is 3/10.

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

ah yes 3/5 is correct ...

OpenStudy (cwrw238):

- for sin x